# How Do LLMs Use Their Depth?

### v1

@@ 4.3 Case Study III: Depth Use during Downstream Tasks
Dataset and Methodology: In this case study, we evaluate the internal prediction dynamics when the model performs four option-constrained downstream tasks - MMLU (Hendrycks et al., 2020), Sentiment Analysis (SST) (Socher et al., 2013), Natural Language Inference (NLI) (Dagan et al., 2005) and Paraphrase Dectection (MRPC) (Dolan & Brockett, 2005). We use a 4-shot set up, where the model is given four examples and is then asked to generate the answer. We then track the mean rank of the generated answer through the different layers of the model. We plot the examples for Pythia-6.9B when answering MMLU and sentiment analysis questions. Plots for other tasks and models can be found in Appendix C.1.
@@
Results: The results can be seen in Figure 5. In these plots, we track the mean rank of the option choices through the different layers. We see a stark, two-step prediction pattern. In the early layers, the model identifies the valid option choices and gathers them within the top ranks of intermediate layer logits. As a result, the early layers see massive decrease in the ranks of valid option choices for MMLU and SST datasets respectively. This step usually happens within the first half of the model. After the valid options are collected as top-ranked choices, the model reasons through the options in the remainder of the layers. This can be seen within the zoomed-in figures on the top-right corners of Figures 5(a) and 5(b), where we zoom-in on the y-axis. For the MMLU task, the model constantly switches its top prediction, with some mid-layer bias for option A and late-layer bias for options C and D. For the sentiment task, the model defaults to “positive” as the label of choice all the way until the last layer. Similar option biases can be seen in other models (Appendix C.1), with extreme biases in smaller models (GPT2-XL) that reduce with model size. Thus, for downstream tasks, the refinement process happens in the second half of the model continues until the very last layer on average.
@@
Overall, we see that the model uses its depth intelligently by first dividing the downstream task into two steps - collecting the valid option choices in the top-ranks and then reasoning through the options. In accordance with the complexity-aware usage of depth, the model performs the easier of the two tasks of collecting the valid option in the early layers, while it leaves reasoning to deeper layers.
@@ 5 Verifying the validity of TunedLens predictions
The results presented in this paper rely on the predictions generated by TunedLens (Belrose et al., 2023). Thus, it is important to demonstrate that our conclusions reflect the internal representations of LLMs themselves rather than the biases introduced by TunedLens. We consider the possibility that the bias toward high frequency tokens was introduced by TunedLens itself and test it in two ways. First, we compare the probability of high-frequency tokens in each layer of TunedLens as compared to the final LLM layer and find them to be comparable. Second, we train our custom TunedLens by varying update frequencies of high-frequency tokens. We find that, even after reducing the update frequency of the most frequent token of a corpus by a factor of 1000, it still appears with an overwhelmingly high frequency in early layer top-ranked predictions. These tests show that early layer predictions in LLMs being dominated by high-frequency tokens is not a consequence of probe bias but rather a reflection of the information content in the early layer representations. Experimental details and additional experiments on the analysis of TunedLens faitfulness, as well as a comparison with LogitLens, can be found in Appendix D.
@@ 6 Related Work
The LogitLens framework (Nostalgebraist, 2020) allows us to interpret intermediate representations in the token space. Building on this, TunedLens (Belrose et al., 2023) trains lightweight affine probes to make intermediate predictions more faithful, especially for earlier layers, and explicitly frames transformers as performing iterative inference across depth. DecoderLens (Langedijk et al., 2023) extends the lens idea to encoder–decoder models and reports that specific subtasks are handled at lower or mid layers. Our work uses the TunedLens probe to explain the prediction dynamics happening during inferece using a “guess-then-refine” framework and an intelligent use of model depth. The idea of guess refinement was mentioned briefly in the original LogitLens blogpost (Nostalgebraist, 2020); however, that work provided no quantitative evidence of the refinement process, which is what we strive to do here.
@@
Subsequent works to LogitLens also formalize saturation events, showing that top-1 token get locked-in before the last layer (Geva et al., 2022), and later layers continue to adjust margins and competitors, which was later generalized into a sequential lock-in process for top-k tokens (Lioubashevski et al., 2024). Not all forward passes lead to saturation events and our work on the other hand focuses on pre-saturation events and explaining internal prediction dynamics during that regime.
Factual Recall and Lookahead Computations:
@@
Prior work shows that factual knowledge is stored within the MLP layers of LLMs (Geva et al., 2020; 2023; Meng et al., 2022) and also present evidence of lookahead planning in LLMs (Pal et al., 2023; Hernandez et al., 2023; Wu et al., 2024; Jenner et al., 2024; Men et al., 2024a). In our work, we tie these observations with depth-usage while performing multi-token fact recall. We show that the second and third token predictions of a multi-token fact appears much earlier into a model and hence requires less computation, an effect that may be partially explained by lookahead planning for future tokens during the first token prediction.
Internal Mechanisms during In-Context Learning:
@@
Many prior works have studied the internal mechanisms of in-context learning in LLMs (Xie et al., 2021; Dai et al., 2022; Von Oswald et al., 2023), but do not focus on how the ranks of labels get modified through the forward pass. Closest to our work is Wang et al. (2023) which shows that shallow layers store task semantics in label tokens, which is retrieved by later layers. Our work presents an orthogonal explanation of this process from the point of view of prediction dynamics and shows that shallow layers elevate the label tokens as top-ranked tokens, while deeper layers reason between them to perform the final selection.
Dynamic-Depth Models and Early-Exiting:
Many recent (Gromov et al., 2024; Men et al., 2024b; Fan et al., 2024) and prior works (Bolukbasi et al., 2017; Huang et al., 2017; Elbayad et al., 2019) have shown that the entire depth of the model is not required for generation and have used this phenomenon to try and improve inference efficiency. Our work provides evidence that early exiting conflicts with the LLM’s natural “guess-then-refine” prediction dynamics. Early exiting strategies may choose to exit while refinement is still ongoing, thus increasing the token selection error rate and over-predicting
@@ 7 Conclusion
In this paper, we explain the internal prediction dynamics of LLMs during inference by proposing a “Guess-then-Refine” framework. We show that early LLM layers promote high-frequency tokens to top-ranked predictions in early layers. These high-frequency tokens act as statistical guesses made by LLMs due to lack of enough contextual information about the input in the early layers. As further processing happens, these early layer guesses undergo massive refinement to eventually yield contextually appropriate tokens. We also show that LLMs leverage their depth flexibly in a task-dependent manner, by using early layers to perform easier tasks like predicting function tokens or identifying valid response options in a constrained-choice task setup, while using later layers for more complex processing like predicting content words, recalling facts, and reasoning. Together, our findings show that LLMs are early statistical guessers and late contextual integrators that use their depth flexibly based on the complexity of predicted token or subtask.
Acknowledgements
This work was supported in part by the NVIDIA Academic Grant Program award. We thank LITlab members for their feedback on this work.
@@ C.1 Downstream Task Ablations
MMLU (GPT2-XL)
MRPC (GPT2-XL)
NLI (GPT2-XL)
SST (GPT2-XL)
MMLU (Llama-2-7B)
MRPC (Llama-2-7B)
NLI (Llama-2-7B)
SST (Llama-2-7B)
MMLU (Llama-3-8B)
MRPC (Llama-3-8B)
NLI (Llama-3-8B)
SST (Llama-3-8B)
MMLU (Pythia-6.9B)
MRPC (Pythia-6.9B)
NLI (Pythia-6.9B)
SST (Pythia-6.9B)
Figure 10: Mean rank by choice for each model–task pair. Each row corresponds to a model and each column to a downstream task.
MMLU (GPT2-XL, zoomed)
MRPC (GPT2-XL, zoomed)
NLI (GPT2-XL, zoomed)
SST (GPT2-XL, zoomed)
MMLU (Llama-2-7B, zoomed)
MRPC (Llama-2-7B, zoomed)
NLI (Llama-2-7B, zoomed)
SST (Llama-2-7B, zoomed)
MMLU (Llama-3-8B, zoomed)
MRPC (Llama-3-8B, zoomed)
NLI (Llama-3-8B, zoomed)
SST (Llama-3-8B, zoomed)
MMLU (Pythia-6.9B, zoomed)
MRPC (Pythia-6.9B, zoomed)
NLI (Pythia-6.9B, zoomed)
SST (Pythia-6.9B, zoomed)
Figure 11: Zoomed-in mean rank by choice for each model–task pair. Each row corresponds to a model and each column to a downstream task.

### v2

@@ 4.1 Case Study I: Depth Use during Downstream Tasks
Dataset and Methodology: In this case study, we evaluate the internal prediction dynamics when the model performs four fixed-option downstream tasks - MMLU (Hendrycks et al., 2020), Sentiment Analysis (SST) (Socher et al., 2013), Natural Language Inference (NLI) (Dagan et al., 2005) and Paraphrase Dectection (MRPC) (Dolan & Brockett, 2005). We use a 4-shot set up, where the model is given four examples and is then asked to generate the answer. We then track the mean rank of the generated answer through the different layers of the model, where the intermediate logits are created using TunedLens.
@@
Results: The results can be seen in Figure 4. We see a clear two-step prediction pattern. In the early layers, the model identifies the valid option choices and gathers them within the top ranks of intermediate layer logits (option-collection). As a result, the early layers see massive decrease in the ranks of valid option choices for MMLU and SST datasets respectively. This step usually happens within the first half of the model. After the valid options are collected as top-ranked choices, the model reasons through the options in the remainder of the layers (option-reasoning). This can be seen within the zoomed-in figures on the top-right corners of Figures 4(a) and 4(b), where we zoom-in on the y-axis. For the MMLU task, the model constantly switches its top prediction, with some mid-layer bias for option A. For the sentiment task, the model defaults to “positive” as the label of choice all the way until the last layer on average. Similar option biases can be seen in other models (Appendix C.1), with extreme biases in smaller models (GPT2-XL).
@@
Overall, we see that the model acts in accordance with complexity-driven use principle, and performs the easier of the two tasks of collecting the valid option in the early layers, while it leaves reasoning to deeper layers.
4.1.1 Causal Intervention by Activation Patching
Setup: We perform activation patching experiments (Meng et al., 2022) to validate our findings. We consider two tasks, MMLU and SST, where SST is the source task and MMLU is the target task. We take the activation vector corresponding to the generated token position when processing the source task (SST) at layer $l$ and patch it during the forward pass of the target task (MMLU) at the same layer. We then look at the response generated during the target task (MMLU). The expectation is that if there are two distinct phases of option-collection and option-reasoning, then transfer of activation vectors during the option-collection phase should still give the model time to collect valid options, whereas transfer of activation vector during option-reasoning phase should generate one of the two options belonging to the SST categories even when geenrating the answer to an MMLU question.
(a) Pythia-6.9B (Fact)
(b) Llama3-8B (Fact)
(c) Pythia-6.9B (POS)
(d) Llama3-8B (POS)
Figure 5: Earliest crossing thresholds of predicted tokens for multi-token facts (top) and POS (bottom). We determine which token the model predicts at the final layer and then track the earliest layer (x axis) at which the TunedLens rank of that token crosses a given threshold (y axis). (a, b) Case Study II: We plot tokenwise results for fact recall responses that consist of 1 token, 2 tokens, or 3 tokens. E.g., 3-token facts consist of 3 tokens at positions (pos) 1, 2, and 3. In both models, the first tokens of multi-token facts appear much later than subsequent tokens. (c, d) Case Study III: When predicted tokens are grouped by POS category, tokens corresponding to function words (DET, ADP) and punctuation appear much earlier than content words (ADJ, VERB, NOUN).
Results: The results are shown in Figure 4(c), where the model answers an MMLU question and receives an activation patch from an SST question. The y-axis of the figure measures percentage of times the predicted answer belongs to either one of the MMLU options (green line), or one of the SST options (blue line), or neither (yellow line). This plot shows replacement of activations for Pythia 6.9B. We see that if we replace the activation from SST to MMLU processing in early layers, which is the option-collection phase, the model is still able to collect the MMLU options to top ranks and produce one of the MMLU options as the final answer. However, if this transfer happens beyond the transition layers, which is the option-reasoning phase, the output belongs to one of the SST options. This experiment clearly demonstrates the presence of a two-phase answering algorithm followed by the model. Note that this activation replacement experiment is done independent of the TunedLens probe. Results for other models show similar outcomes as shown in Figure 19.
We also define phase transition layers between the two phases using the TunedLens probe, which is the layer at which the model transitions from option-collection to option-reasoning mode. We define transition at the earliest layer at which all the top ranked tokens choices using the TunedLens probe belong to the available options for the task. This means that for MMLU, the top four ranked tokens should belong to each of the four option choices in MMLU (A/B/C/D). Similarly, for SST, the top two ranked tokens should belong two each of the two options of the task (positive/negative). We see that the average transition layer when calculated using TunedLens matches very nicely with the transition happening in Figure 4(c) (Appendix), which is measured independent of the probe. This shows that intermediate logits generated using TunedLens are predictive of actual model behavior.
4.2 Case Study II: Depth Use during Multi-Token Fact Recall
Methodology: In this case study, we track the predicted tokens through intermediate layers during fact recall using the MQuAKE dataset (Zhong et al., 2023), which contains a variety of single-token and multi-token facts. We split the dataset into three different cases where the answer contains one, two or three tokens and look at how the rank of the predicted tokens proceed for each generated token of the answer. More information about the dataset statistics can be found in Appendix C.
An example query in this setting is - “The Statue of Liberty is located in” and the expected generated tokens are “New York City” (Figure 1 (c)). For each of the output tokens in the answer, we use TunedLens to trace the rank of the output token at an intermediate layer. The earliest layer at which the predicted token crosses a particular rank threshold is recorded. We record the earliest crossing to track the minimum computation required to generate a token. Additionally, we present internal computations only in scenarios where the model predicts the correct answer. We explain the reasons behind this and a comparison for scenarios where model also predicts incorrect answers in Appendix H. We also measure the layer at which predicted token crosses a given rank threshold for the final time and find similar qualitative results, as shown in Appendix E.
(a) Pythia 6.9B POS
(b) Pythia 6.9B 2-Token Fact
(c) Pythia 6.9B 3-Token Fact
Figure 6: Early exiting experiments. To test our hypothesis that easy to predict token categories finish processing earlier more often, we find the Top-1 match rate for an intermediate-layer decoded token with the final prediction - the match is successful if the Top-1 ranked prediction at a given layer matches the final prediction. The decoding is done via TunedLens. In (a), we see that the match rate in early layers is much larger for function tokens belonging to DET, ADP and PUNCT tokens, whereas it is much lower for content token categories of ADJ, VERB and NOUN. In (b, c), we see that once the first token is predicted, the subsequent tokens are easier to predict at earlier layers. The Top-1 match rate for the second token in 2-token facts, and the third token in 3-token facts is much higher in earlier layers. Similar results can be seen for other models in Figure 18.
Results: Figures 5(a) and 5(b) show the earliest layer at which the correct fact recall token crosses a given rank threshold. We see that the first appearance of the single-token answer happens much earlier into a model (layer 15 for Pythia-6.9B, layer  20 for Llama3-8B) compared to the first token of the multi-token facts (layer 25 for both models). This shows that recalling multi-token facts, especially the first token of a multi-token fact, is harder for the model compared to a single token fact. Predicting a single token fact only requires accurate construction of a single token, while predicting a multi-token fact also requires a model to anticipate the future tokens that complete fact recall accurately, which may make the recalling multi-token facts harder. Also, we see that all computations for fact recall tokens progress much slower ( layer 15) than function word categories from Case Study III ( layer 5). This means that recalling facts is not as easy as generating DET or ADP categories and requires much larger depth.
Critically, while computing a multi-token fact, the earliest appearance of the first token in the multi-token fact requires the largest number of layers, but the subsequent tokens become top-ranked much sooner. For instance, for a 3-token fact, the first token on average emerges around layer 27 for Pythia-6.9B, whereas the subsequent second and third tokens emerge around layers 20 and 12 respectively. This pattern represents another example of a graded use of model depth, where harder to do tasks, like predicting the first token in a multi-token fact, requires additional depth, while subsequent tokens that are conditionally easier to predict given the first token, appear at shallower depths.
Early-Decoding Experiments: To validate the information processing trajectories during multi-token fact recall, we perform early exiting experiments and calulate Top-1 match rate at intermediate layers. We do this using TunedLens and calculated the percentage of times the intermediate Top-1 rank prediction matches the final prediction at an intermediate layer. These experiments have been used in prior work to validate intermediate computations performed by the model (Din et al., 2023; Geva et al., 2022, 2020). The results can be seen in Figures 6(b) and 6(c), where the Top-1 match rate for subsequent tokens is much higher in early layers. This effect is especially stark when observing the second token in a 2-token fact and the third token in a 3-token fact for Pythia 6.9B.
4.3 Case Study III: Depth Use by POS Category
Methodology: Finally, we study the depth use strategy in LLMs during next-token prediction by tracking how the predicted token becomes top-ranked through the model depth similar to section 4.2. The predicted token is classified into one of the following part-of-speech (POS) categories - determiners (DET), adpositions (ADP), punctuations (PUNCT), adjectives (ADJ), verbs, and nouns. We use POS to categorize generated tokens because they give us an indication of the role a generated word plays in a sentence. POS categories like DET and ADP contain function words that do not provide much meaning to a sentence, whereas other categories like adjectives, verbs, and nouns are made up of content words where the main information content of a sentence is present.
For example, if the input prefix sentence is “Today is a sunny” and the model predicts “day” as the generation, we first use Spacy to classify the POS of the predicted word. Then, we use TunedLens to trace the predicted rank of “day” at different layers within the model. The first layer inside a model at which the predicted token crosses a particular rank threshold is recorded. We plot the average layer at which a particular rank threshold is crossed by the predicted token at intermediate layers.
Results: Figures 5(c) and 5(d) show the average number of layers required for a predicted token to first cross a rank-threshold. Function words (DET and ADP) and punctuation (PUNCT) progress and reach rank-1 much faster than content words (ADJ, VERB, NOUN). For example, when the predicted word is a DET, it first reaches rank 1 on average around layer 5 for both Pythia-6.9B and Llama3-8B, while other content words categories reach rank-1 much deeper into a model, closer to layer 20. The same phenomenon is observed for the final crossing metric as shown in Figure 16 (Appendix). These results are also consistent with the finding in Section 3: DET, ADP, and PUNCT contain primarily high-frequency tokens and are therefore predicted early in the model, at the initial guessing stage. Content words, however, must rely on contextual inference in middle layers to be guessed correctly.
Early-Decoding Experiments: To validate that the information processing trajectories when predicting function tokens proceeds faster than content tokens, we perform early decoding experiments similar to section 4.2. The results are present in Figure 6(a), where we clearly see that the Top-1 match rate in early layers is much higher for function tokens (DET, ADP) and PUNCT, whereas it is much lower for content token categories. To quantify this difference, at layer 10, it is almost twice as likely for the Top-1 prediction to be the final token when the final token belongs to the function token or PUNCT categories than when it belongs to content token categories. These results have direct implications for future work in early-exiting.
@@ 5 Verifying the validity of TunedLens predictions
In this section, we summarize additional experiments used to evaluate the possibility of a frequency bias introduced by TunedLens. First, we compare the probability of high-frequency tokens in each layer of TunedLens as compared to the final LLM layer and find them to be comparable. Second, we train our custom TunedLens by varying update frequencies of high-frequency tokens. We find that, even after reducing the update frequency of the most frequent token of a corpus by a factor of 1000, it still appears with an overwhelmingly high frequency in early layer top-ranked predictions. These tests show that early layer predictions in LLMs being dominated by high-frequency tokens is not a consequence of probe bias but rather a reflection of the information content in the early layer representations. Experimental details and additional experiments on the analysis of TunedLens faitfulness, as well as a comparison with LogitLens, can be found in Appendix D.
@@ 6 Related Work
Layerwise Decoding and Iterative Refinement: The LogitLens framework (Nostalgebraist, 2020) allows us to interpret intermediate representations in the token space. Building on this, TunedLens (Belrose et al., 2023) trains lightweight affine probes to make intermediate predictions more faithful, especially for earlier layers. DecoderLens (Langedijk et al., 2023) extends the lens idea to encoder–decoder models. Our work uses the TunedLens probe to explain the prediction dynamics during inference using a “guess-then-refine” framework and an dynamic use of model depth. The idea of guess refinement was briefly mentioned in the original LogitLens blogpost (Nostalgebraist, 2020); however, that work provided no quantitative evidence of the refinement process, which is what we strive to do here.
@@
Subsequent works to LogitLens also formalize saturation events, showing that top-1 token get locked-in before the last layer (Geva et al., 2022), and later layers continue to adjust margins and competitors, which was later generalized into a sequential lock-in process for top-k tokens (Lioubashevski et al., 2024). Our work on the other hand focuses on pre-saturation events and explaining internal prediction dynamics during that regime.
@@
Fact Recall and Lookahead Computations: Prior work shows that factual knowledge is stored within the MLP layers (Geva et al., 2020, 2023; Meng et al., 2022) and also present evidence of lookahead planning in LLMs (Pal et al., 2023; Hernandez et al., 2023; Wu et al., 2024; Jenner et al., 2024; Men et al., 2024a). In our work, we tie these observations with depth-usage while performing multi-token fact recall by showing that the subsequent token predictions of a multi-token fact require less computation. Merullo et al. (2024) study internal dynamics during fact recall and show early layers generate function arguments to retrieve factual information, which is very different from our work.
@@
Internal Mechanisms during In-Context Learning: Many prior works have studied the internal mechanisms of in-context learning in LLMs (Xie et al., 2021; Dai et al., 2022; Von Oswald et al., 2023), but do not focus on how the ranks of labels get modified through the forward pass. Lepori et al. (2025) explores the idea of refinement during the forward pass focusing on contextualization errors. Closest to our work is Wang et al. (2023), who shows that shallow layers store task semantics in label tokens, which is retrieved by later layers. Our work presents an orthogonal explanation of this process from the point of view of prediction dynamics and shows that shallow layers elevate the label tokens as top-ranked tokens, while deeper layers reason between them to perform the final selection.
Dynamic-Depth Models and Early-Exiting: Many recent (Gromov et al., 2024; Men et al., 2024b; Fan et al., 2024) and prior works (Bolukbasi et al., 2017; Huang et al., 2017; Elbayad et al., 2019) have shown that the entire depth of the model is not required for generation and have used this phenomenon to try and improve inference efficiency. Csordás et al. (2025) build on these studies and show that models with larger number of layers do not neccesarily incorporate newer but rather perform fine-grained adjustments. Our work on the other hand tries to provide a structure to the usage of the depth given to LLMs and explain it using the proposed guess-then-refine framework.
@@ 7 Conclusion
In this paper, we explain the internal prediction dynamics of LLMs during inference by proposing a “Guess-then-Refine” framework. We show that early LLM layers promote high-frequency tokens to top-ranked predictions in early layers. These high-frequency tokens act as statistical guesses made by LLMs due to lack of enough contextual information about the input in the early layers. As further processing happens, these early layer guesses undergo massive refinement to eventually yield contextually appropriate tokens. We also show that LLMs leverage their depth flexibly in a task-dependent manner, by using early layers to perform easier tasks like predicting function tokens or identifying valid response options, while using later layers for more complex processing like predicting content words, recalling facts, and reasoning. Together, our findings show that LLMs are early statistical guessers and late contextual integrators that use their depth flexibly based on the complexity of predicted token or subtask. Some avenues of future work includes a similar analysis for reasoning and chain-of-thought tasks and more recent reasoning models.
@@ C.1 Downstream Task Ablations
(a) GPT2-XL — MMLU
(b) GPT2-XL — MRPC
(c) GPT2-XL — NLI
(d) GPT2-XL — SST
(e) Llama-2-7B — MMLU
(f) Llama-2-7B — MRPC
(g) Llama-2-7B — NLI
(h) Llama-2-7B — SST
(i) Llama-3-8B — MMLU
(j) Llama-3-8B — MRPC
(k) Llama-3-8B — NLI
(l) Llama-3-8B — SST
(m) Pythia-6.9B — MMLU
(n) Pythia-6.9B — MRPC
(o) Pythia-6.9B — NLI
(p) Pythia-6.9B — SST
Figure 11: Mean rank by choice across datasets for multiple models
(a) GPT2-XL — MMLU (zoomed)
(b) GPT2-XL — MRPC (zoomed)
(c) GPT2-XL — NLI (zoomed)
(d) GPT2-XL — SST (zoomed)
(e) Llama-2-7B — MMLU (zoomed)
(f) Llama-2-7B — MRPC (zoomed)
(g) Llama-2-7B — NLI (zoomed)
(h) Llama-2-7B — SST (zoomed)
(i) Llama-3-8B — MMLU (zoomed)
(j) Llama-3-8B — MRPC (zoomed)
(k) Llama-3-8B — NLI (zoomed)
(l) Llama-3-8B — SST (zoomed)
(m) Pythia-6.9B — MMLU (zoomed)
(n) Pythia-6.9B — MRPC (zoomed)
(o) Pythia-6.9B — NLI (zoomed)
(p) Pythia-6.9B — SST (zoomed)
Figure 12: Mean rank by choice across datasets for multiple models (zoomed-in views